A body vibrate in simple harmonic motion with a frequency of 50hz and an amplitude of 4cm. Find



1.period



2. Acceleration at the middle and at the end of the path of oscillation

1
Expert's answer
2022-03-18T09:15:29-0400

1.

The period of oscillations

"T=\frac{1}{f}=\frac{1}{50\rm\: Hz}=0.02\:\rm s"

2.

The acceleration at the middle point is zero.

The acceleration at the end point is equal

"a_{\max}=A\omega^2=4\pi^2Af^2\\
=4\times 3.14^2\times0.04\;\rm m\times (50\: Hz)^2=3944\: m/s^2"